OpenStudy (anonymous):

Graph the equation by plotting 3 points x - y =1

OpenStudy (anonymous):

probably easier if you write this as \[y=x-1\]then just pick any x, substitute to get y for example \[x=0\] then \[y = 0-1=-1\] so you get (0,-1) and if \[x=2\] you get \[y=2-1=1\] so plot (2,1) and so on

OpenStudy (anonymous):

can you please help me with another???

OpenStudy (anonymous):

sure name it

OpenStudy (anonymous):

ok hold please

OpenStudy (anonymous):

are you good at math and linear equations?

OpenStudy (anonymous):

name one we can try it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok graph the linear equation, find three points that solve the equation -x - 3y= -6

OpenStudy (anonymous):

ok lets pick an x i pick 0 so first coordinate is 0 we solve for y \[0-3y=-6\] \[-3y=-6\] \[y=\frac{-6}{-3}=2\] giving a point (0,2)

OpenStudy (anonymous):

now you pick a y. (hint, pick y = 0)

OpenStudy (anonymous):

so it would be 0,3

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

well actually since you picked y = 0 it would be \[-x-0=-6\] \[-x=-6\] \[x=6\] and the point would be (6,0)

OpenStudy (anonymous):

btw if i get (0,2) you cannot also get (0,3). the first coordinates have to be different.

OpenStudy (anonymous):

so the 2 points are 6,0 0,2 how do i get the 3rd one

OpenStudy (anonymous):

?????? help?

OpenStudy (anonymous):

pick some other number for x and solve for y. for example if you pick x = 3 you get \[-3-3y=-6\] \[-3y=-6+3\] \[-3y=-3\] \[y=\frac{-3}{-3}\] \[y=1\] so you have the third point is (3,1)